OpenStudy (anonymous):

Will Fan and medal!!! Find an equation of a parabola with a vertex at the origin and directrix at y=-3.5

OpenStudy (anonymous):

A. y=-1/14x^2 B. y=1/14x^2 C. y=1/14y^2 D. y=-1/14y^2

OpenStudy (anonymous):

The answer is B for anyone else who needs it